# Zero-Shot Sim-to-Real Visual Quadrotor Control with Hard Constraints

### v1

@@ Abstract
We present the first framework demonstrating zero-shot sim-to-real transfer of visual control policies learned in a Neural Radiance Field (NeRF) environment for quadrotors to fly through racing gates. Robust transfer from simulation to real flight poses a major challenge, as standard simulators often lack sufficient visual fidelity. To address this, we construct a photorealistic simulation environment of quadrotor racing tracks, called FalconGym, which provides effectively unlimited synthetic images for training. Within FalconGym, we develop a pipelined approach for crossing gates that combines (i) a Neural Pose Estimator (NPE) coupled with a Kalman filter to reliably infer quadrotor poses from single-frame RGB images and IMU data, and (ii) a self-attention-based multi-modal controller that adaptively integrates visual features and pose estimation. This multi-modal design compensates for perception noise and intermittent gate visibility. We train this controller purely in FalconGym with imitation learning and deploy the resulting policy to real hardware with no additional fine-tuning. Simulation experiments on three distinct tracks (circle, U-turn and figure-8) demonstrate that our controller outperforms a vision-only state-of-the-art baseline in both success rate and gate-crossing accuracy. In 30 live hardware flights spanning three tracks and 120 gates, our controller achieves a 95.8% success rate and an average error of just 10 cm when flying through 38 cm-radius gates.
@@ I INTRODUCTION
Figure 1: Our simulation-trained autonomous quadrotor, equipped with an ArduCam RGB camera and an IMU-enabled Navio2 on a Raspberry Pi 3, can successfully fly through a sequence of small 38 cm-radius gates with an average of only 10 cm error to gates center.
Recent advances in neural scene representations, such as Neural Radiance Fields (NeRF) [3] and Gaussian Splatting [4], provide an opportunity for creating photorealistic simulation environments, which can be used for training visual control policies. Although prior works have shown promising zero-shot neural-scene to real transfer for humanoid robots [5] and grasping [6], few have focused on quadrotors, which demand higher speeds and control frequencies than ground robots. SOUS VIDE [7] is a notable exception, which uses Gaussian Splat [4] for neural scene representation and achieves impressive zero-shot drone navigation by distilling an expert MPC controller to an lightweight end-to-end policy.
@@
In summary, our contribution is threefold. First, we demonstrate the first framework to achieve zero-shot sim-to-real transfer of NeRF-trained visual control policies for quadrotors navigating through a sequence of small gates. Second, we develop FalconGym, a photorealistic simulation environment for agile, vision-based quadrotor flight, providing an open benchmark for visual flight control research. Third, we make two innovations that make our neural controller’s architecture practical:(i) A one-shot Neural Pose Estimator (NPE) coupled with a Kalman filter; (ii) A self-attention-based sensor fusion scheme that integrates vision features and pose estimation for quadrotor flight.
@@ II RELATED WORK
Neural scene representations like NeRF [3] and Gaussian Splat [4] have recently emerged as powerful tools for photorealistic 3D scene reconstruction. Variants and extensions have been applied to diverse robotics tasks, including 3D scene editing [8], pose estimation [9] and navigation [10]. NeRF2Real [5] and RialTo [6] demonstrate the potential sim-to-real transfer for ground robots and robot arm manipulation. [7] achieves zero-shot drone navigation using policies trained in Gaussian splat scenes. In addition to using NeRF representations for training, our work also uses a completely different solution architecture than [7] and we solve the navigation problem with harder constraints in the shape of 38 cm-radius gates.
@@
Sim-to-real transfer has long been a cornerstone of robotics research, valued for its efficient and risk-free training. Traditional methods such as domain randomization [11] and system identification [12] are commonly employed to narrow the gap between simulation and reality. These techniques and their variations have enabled successful sim-to-real deployments in quadruped locomotion [13], bipedal walking [14], robotic grasping [15], and autonomous driving [16].
@@
Figure 2: Our FalconGym can render photorealistic images from different poses. The left column shows four real world images captured by the Arducam in our flight area, while the right column displays the corresponding images rendered by FalconGym at the same coordinates.
@@ III METHODOLOGY
Figure 3: The architecture of our visual flight system consists of a perception module and a control module. In the blue box, the perception module processes onboard camera images using a ViT-based Neural Pose Estimator (NPE) and improves its estimates with IMU data via a Kalman filter. The purple box details the control module, where a self-attention mechanism fuses state embeddings with vision features to generate control actions.
@@ III-C Controller Architecture: Self-Attention
Figure 5: Qualitative evaluation of the perception module $P$ is shown. Blue dots represent a subset of $\mathcal{D}_{Q}$ (see Section III-D), which compare the pose estimation from perception module $P$ against the ground truth pose. The black line indicates perfect estimation, while the orange shaded area denotes the 10–90% quantile range of perception errors, which is used to augment the controller training dataset $\mathcal{D}_{C}$. The close clustering of the blue dots around the black line demonstrates that our perception module achieves relatively accurate pose estimation.
@@ III-D2 Quantile-Based Perception Error Modeling
Inspired by the “perception contract” concept [34, 35], we explicitly model the error between ground-truth and estimated states. First, we employ the expert controller (with domain randomized control) in FalconGym, to collect a dataset $\mathcal{D}_{Q}=\{(\hat{p}_{t},p_{t})\}_{t=1}^{K}$ where $\hat{p}_{t}=P(F(p_{t}),z_{t})$ is the pose estimation by NPE and Kalman filter (Section III-B). We gathered $K\sim 10$k over 500 runs in FalconGym. Next, we apply quantile regresion [36] to $\mathcal{D}_{Q}$ to estimate the conditional 10th and 90th percentiles of the estimation error distribution, denoted $Q_{0.1}(p)$ and $Q_{0.9}(p)$. These quantiles define an empirical 80% prediction interval for the error term $\hat{p}-p$ such that
@@ III-D4 Multi-Modal Controller Learning
TABLE I: Evaluation of the perception module. The table compares the Neural Pose Estimator (NPE) used alone versus with Kalman filtering, and evaluates perception module’s performance of zero-shot sim-to-real transfer
@@
Track Pose Error [cm] Yaw Error [deg] Pose Error [cm] Yaw Error [deg] Pose Error [cm] Yaw Error [deg] Pose Error [cm] Yaw Error [deg]
@@ IV-A Closed-Loop System Setup in Hardware
We conduct our hardware experiments on a quadrotor measuring 30cm $\times$ 34cm $\times$ 11cm, equipped with a Raspberry Pi 3® and a Navio2® board for flight control, as shown in Figure 1. The Navio2® incorporates an onboard IMU and streams acceleration data at up to 50hz. An ArduCam global-shutter camera is also mounted on the quadrotor, in order to reduce both computation and communication overhead, we operate the camera at 20Hz and downsample the image size to $320\times 240$.
All sensor data, including camera images and IMU measurements, are published to ROS topics on the Pi. An offboard workstation, connected on the same subnet as the Pi and equipped with an NVIDIA RTX 4090 GPU, subscribes to these topics. And then we perform the perception (Section III-B) and controller inference (Section III-C) offboard, and finally sends the computed control actions back to the Pi via another ROS topic. Empirically, we measure an end-to-end latency of about 39ms between when sensor data are sent and when control commands are received. This allows us to maintain a stable 20Hz (50ms delay) control loop in the hardware. To be consistent, we also use a $\Delta t=0.05s$ step size (i.e. 20Hz) for training in FalconGym as well.
@@ IV-D Controller Evaluation
Following the visualzied gate-crossing trajectories shown in Figure 4, we next present a quantitative analysis that addresses two main questions:(i) Does our Multi-Modal (MM) controller outperform a vision-only state-of-the-art (SoTA) baseline? (ii) Can our MM approach transfer effectively from FalconGym to real hardware?
@@
TABLE III: Evaluation of four controllers based on Success Rate (SR) and Mean Gate Error (MGE) in FalconGym and real world.
@@
Method SR $\uparrow$ MGE (cm) $\downarrow$ SR $\uparrow$ MGE (cm) $\downarrow$
@@ V CONCLUSION
In this work, we present a novel visual control framework for quadrotor gate navigation that achieves zero-shot Sim2Real transfer. Leveraging our photorealistic FalconGym simulation environment, we train a visual controller that integrates visual features and pose estimation through a self-attention-based sensor fusion mechanism, making it robust to perception errors. Experiments in both FalconGym and hardware demonstrates high performance of our approach.
Despite these promising results, several factors may lead to failures: a substantial visual gap between FalconGym and the real world; discrepancies between the simulated dynamics in FalconGym and the true quadrotor; or significant offboard computation and communication delays could compromise our approach’s effectiveness. Future work can pursue several directions. First, latest neural rendering techniques such as Gaussian Splatting [39] could replace our current NeRF-based scene representation, providing faster rendering and higher image fidelity. Second, we plan to extend our simulation beyond the simplistic double-integrator dynamics, incorporating more realistic physics-based models [40] to better support high-speed or aggressive maneuvers. Finally, to reduce reliance on offboard computing, we aim to distill both the perception and control modules into more compact models that can run entirely onboard.

### v2

# FalconGym: A Photorealistic Simulation Framework for Zero-Shot Sim-to-Real Vision-Based Quadrotor Navigation

@@ Abstract
We present a novel framework demonstrating zero-shot sim-to-real transfer of visual control policies learned in a Neural Radiance Field (NeRF) environment for quadrotors to fly through racing gates. Robust transfer from simulation to real flight poses a major challenge, as standard simulators often lack sufficient visual fidelity. To address this, we construct a photorealistic simulation environment of quadrotor racing tracks, called FalconGym, which provides effectively unlimited synthetic images for training. Within FalconGym, we develop a pipelined approach for crossing gates that combines (i) a Neural Pose Estimator (NPE) coupled with a Kalman filter to reliably infer quadrotor poses from single-frame RGB images and IMU data, and (ii) a self-attention-based multi-modal controller that adaptively integrates visual features and pose estimation. This multi-modal design compensates for perception noise and intermittent gate visibility. We train this controller purely in FalconGym with imitation learning and deploy the resulting policy to real hardware with no additional fine-tuning. Simulation experiments on three distinct tracks (circle, U-turn and figure-8) demonstrate that our controller outperforms a vision-only state-of-the-art baseline in both success rate and gate-crossing accuracy. In 30 live hardware flights spanning three tracks and 120 gates, our controller achieves a 95.8% success rate and an average error of just 10 cm when flying through 38 cm-radius gates.
@@ I INTRODUCTION
Figure 1: Our autonomous quadrotor, equipped with an ArduCam RGB camera and an IMU-enabled Navio2 on a Raspberry Pi 3, can successfully fly through a sequence of small 38 cm-radius gates with an average of only 10 cm error to gates center.
Recent advances in neural scene representations, such as Neural Radiance Fields (NeRF) [3] and Gaussian Splatting [4], provide an opportunity for creating photorealistic simulation environments, which can be used for training visual control policies. Although prior works have shown promising zero-shot neural-scene to real transfer for humanoid robots [5] and grasping [6], few have focused on quadrotors, which demand higher speeds and control frequencies than ground robots. SOUS VIDE [7] is a notable exception, which uses Gaussian Splat [4] for neural scene representation and achieves impressive zero-shot drone navigation by distilling an expert MPC controller to a lightweight end-to-end policy.
@@
In summary, our contribution is threefold. First, we demonstrate a novel framework to achieve zero-shot sim-to-real transfer of NeRF-trained visual control policies for quadrotors navigating through a sequence of small gates. Second, we develop FalconGym, a photorealistic simulation environment for agile, vision-based quadrotor flight, providing an open benchmark for visual flight control research. Third, we make two innovations that make our neural controller’s architecture practical:(i) A one-shot Neural Pose Estimator (NPE) coupled with a Kalman filter; (ii) A self-attention-based sensor fusion scheme that integrates vision features and pose estimation for quadrotor flight.
@@ II RELATED WORK
Neural scene representations like NeRF [3] and Gaussian Splat [4] have recently emerged as powerful tools for photorealistic 3D scene reconstruction. Variants and extensions have been applied to diverse robotics tasks, including 3D scene editing [8], pose estimation [9], grasping [10] and navigation [11]. NeRF2Real [5] and RialTo [6] demonstrate the potential sim-to-real transfer for ground robots and robot arm manipulation. [7] achieves zero-shot drone navigation using policies trained in Gaussian splat scenes. In addition to using NeRF representations for training, our work also uses a completely different solution architecture than [7] and we solve the navigation problem with harder constraints in the shape of 38 cm-radius gates.
@@
Sim-to-real transfer has long been a cornerstone of robotics research, valued for its efficient and risk-free training. Traditional methods such as domain randomization [12] and system identification [13] are commonly employed to narrow the gap between simulation and reality. These techniques and their variations have enabled successful sim-to-real deployments in quadruped locomotion [14], bipedal walking [15], and autonomous driving [16].
@@
Figure 2: FalconGym can render photorealistic images from different poses. The left column shows four real world images captured by the Arducam in our flight area, while the right column displays images rendered by FalconGym at the same coordinates.
@@ III METHODOLOGY
Figure 3: The architecture of our visual flight system consists of a perception module and a control module. In the blue box, the perception module processes onboard camera images using a ViT-based Neural Pose Estimator (NPE) and improves its estimates with IMU data via a Kalman filter. The purple box details the control module, where a self-attention mechanism fuses state embeddings with vision features.
@@ III-C Controller Architecture: Self-Attention
Figure 5: Qualitative evaluation of the perception module $P$ is shown. Blue dots represent a subset of $\mathcal{D}_{Q}$ (see Section III-D), which compare the pose estimation from perception module $P$ against the ground truth pose. The black line indicates perfect estimation, while the orange shaded area denotes the 10–90% quantile range of perception errors, which is used to augment the controller training dataset $\mathcal{D}_{C}$. The close clustering of tblue dots around the black line demonstrates that our perception module achieves accurate pose estimation.
@@ III-D2 Quantile-Based Perception Error Modeling
Inspired by the “perception contract” concept [34, 35], we explicitly model the error between ground-truth and estimated states. First, we employ the expert controller (with domain randomized control) in FalconGym, to collect a dataset $\mathcal{D}_{Q}=\{(\hat{p}_{t},p_{t})\}_{t=1}^{K}$ where $\hat{p}_{t}=P(F(p_{t}),z_{t})$ is the pose estimation by NPE and Kalman filter (Section III-B). We gathered $K\sim 10$k over 500 runs in FalconGym. Next, we apply quantile regression [36] to $\mathcal{D}_{Q}$ to estimate the conditional 10th and 90th percentiles of the estimation error distribution, denoted $Q_{0.1}(p)$ and $Q_{0.9}(p)$. These quantiles define an empirical 80% prediction interval for the error term $\hat{p}-p$ such that
@@ III-D4 Multi-Modal Controller Learning
TABLE I: Evaluation of the perception module. The table compares the Neural Pose Estimator (NPE) used alone versus with Kalman filtering, and evaluates perception module’s performance of zero-shot sim-to-real transfer.
@@
Track Pose Err. [cm]$\downarrow$ Yaw Err. [deg]$\downarrow$ Pose Err. [cm]$\downarrow$ Yaw Err. [deg]$\downarrow$ Pose Err. [cm]$\downarrow$ Yaw Err. [deg]$\downarrow$ Pose Err. [cm]$\downarrow$ Yaw Err. [deg]$\downarrow$
@@ IV-A Closed-Loop System Setup in Hardware
We conduct our hardware experiments on a quadrotor measuring 30cm $\times$ 34cm $\times$ 11cm, equipped with a Raspberry Pi 3® and a Navio2® board for flight control, as shown in Figure 1. The Navio2® incorporates an onboard IMU and streams acceleration data at up to 50hz. An ArduCam global-shutter camera is also mounted on the quadrotor, in order to reduce both computation and communication overhead, we operate the camera at 20Hz and downsample the image size to 320×240320$\times$ 240320 × 240.
All sensor data, including camera images and IMU measurements, are published to ROS topics on the Pi. An offboard workstation, connected on the same subnet as the Pi and equipped with an NVIDIA RTX 4090 GPU, subscribes to these topics. Then we perform the perception (Section III-B) and controller inference (Section III-C) offboard, and finally sends the computed control actions back to the Pi via another ROS topic. Empirically, we measure an end-to-end latency of about 39ms between when sensor data are sent and when control commands are received. This allows us to maintain a stable 20Hz (50ms delay) control loop in the hardware. To be consistent, we also use a $\Delta t=0.05s$ step size (i.e. 20Hz) for training in FalconGym as well.
@@ IV-D Controller Evaluation
Following the visualized gate-crossing trajectories shown in Figure 4, we next present a quantitative analysis that addresses two main questions:(i) Does our Multi-Modal (MM) controller outperform a vision-only state-of-the-art (SoTA) baseline? (ii) Can our MM approach transfer effectively from FalconGym to real hardware?
@@
TABLE III: Evaluation of controllers by Success Rate (SR) and Mean Gate Error (MGE) in FalconGym and real world.
@@
Method SR $\uparrow$ MGE [cm] $\downarrow$ SR $\uparrow$ MGE [cm] $\downarrow$
@@ V CONCLUSION
In this work, we present a novel visual control framework for quadrotor gate navigation that achieves zero-shot sim-to-real transfer. Leveraging our photorealistic FalconGym simulation environment, we train a visual controller that integrates visual features and pose estimation through a self-attention-based sensor fusion mechanism, making it robust to perception errors. Experiments in both FalconGym and hardware demonstrate high performance of our approach.
Nonetheless, performance can degrade due to visual domain gaps, mismatches between FalconGym’s simulated dynamics and the true quadrotor, and offboard computation or communication delays. Future directions include replacing our NeRF backend with Gaussian Splatting [39] for faster, higher-fidelity rendering; training policies that generalize to unseen tracks; incorporating physics-based dynamics [40] beyond a double-integrator model for aggressive maneuvers; distilling perception and control modules for fully onboard execution; and exploring high-speed flights to handle motion blur and tight perception–action coupling.
ACKNOWLEDGMENT
The authors conducting this research were supported in part by the Air Force Research Laboratories (Award number FA8651-24-1-0007) and the Boeing Company. The authors would like to thank Yangge Li and Ege Yuceel for their valuable feedback.